# Quantum Beatings in Optical Cavities

### v1

@@ I Introduction
Cavity polaritons, quasiparticles arising from the coupling of molecules and photons in an optical cavity, have become a key concept in quantum optics.[1, 2] Along with superconducting circuits, trapped ions, and cold atoms in optical lattices, these hybrid quantum platforms have lent themselves to a diverse set of applications in quantum information, material science and condensed matter physics. Central to their theoretical descriptions are fundamental models describing the interaction between an ensemble of two-level systems (TLS) and the cavity photons, namely, the Tavis-Cummings (TC) model, the Dicke model (DM), and the Pauli-Fierz (PF) model. Common to these models is the prediction of Rabi oscillations, which are widely recognized as the quintessential signature of quantum coherence in the light-matter interaction.[3, 4] In this letter, we predict a more subtle quantum signature that goes beyond the Rabi oscillation in revealing the fine energy structure of light-matter hybrid states (i.e., polaritons) and in differentiating between various theoretical descriptions, even in the weak coupling regime.
@@ II Model Hamiltonians
$H_{PF}=H_{D}+\frac{g^{2}}{\omega_{c}N}\left(J_{x}\right)^{2}$ (3)
@@ III Numerical Analysis
Numerical analysis of the cavity system involves the photon counts, spectral decomposition and energy schema. Simulations were carried out for $N$ from 1 to 15, but we focus primarily on the $N=2$ resonant case for the convenience of the subsequent perturbative analysis. The parameters are $\omega_{c}=\omega_{m}=1.0$, $0.01\leq g\leq 0.12$, and an initial state consisting of a TLS ensemble in the doubly-excited symmetric state and the cavity photon in the vacuum state, $\psi(0)=\ket{s_{2},0}$. To propagate the polariton system, we combine numerical integration of the Schrodinger equation and spectral decomposition [see Eq. (4)].
@@ III.1 Average Photon Count
The average photon count, i.e., the expectation value of photon number operator $a^{{\dagger}}a\right>$, is plotted for the TC, Dicke, and PF models in Fig. 1. These plots illustrate the presence of polariton beatings in the Dicke and PF models but not in the resonant TC model. Additionally, the beating dynamics in the Dicke and PF models are similar–the only noticeable difference between the two occurs between $t=1200$ and $t=1400$, where the oscillation amplitude diminishes. The similarity can be attributed to the DSE term in the PF model, which scales with $g^{2}/N$, deeming it an order of magnitude less significant than the CRW terms. Thus, we will focus on the Dicke model in the subsequent study of the beating dynamics. As shown in SM, a similar beating pattern can be observed on the higher excitation manifold. Further, the variance of the photon count also exhibit a complex beating pattern as shown in SM, suggesting quantum features in photon statistics.[8, 9]
@@ III.3 Energy Schema &Eigenvalue Asymmetry
Evidently, the energy scheme of the TC model is symmetric by construction, so $\alpha=0$ under the RWA in the TC model, but it grows with $g$ for the Dicke model, suggesting that the asymmetry $\alpha$ drives the quantum beating.
@@ V Perturbative analysis of beating
$\Delta E_{\pm}^{(0)}=\frac{\left|\braket{P_{\pm}}{H_{CRW}}{s_{0},0}\right|^{2}}{E_{\pm}}=\frac{g^{2}}{2(2\omega\pm\Omega)}$ (11)
@@
$\Delta E_{i}^{(4)}=\sum_{n=1}^{3}\frac{|\braket{s_{n},4-n}{J_{+}a^{{\dagger}}}{P_{i}}|^{2}}{E_{i}-4\omega}$ (12)
@@
We then define $\Delta E_{i}=\Delta E_{i}^{(0)}+\Delta E_{i}^{(4)}$ since $\Delta E_{0}^{(0)}=0$) such that the perturbed polariton eigenenergies in the Dicke model are
@@
 $\qquad-\,8N\,\cos(\alpha,t)\,\cos(\Omega\,t)\Bigg{]}$ (16)
@@ VI Experimental Relevance
Notably, modern experimental setups are capable of realizing the prediction of polariton beating. Consider a feasible physical system [17] with parameters $\frac{\omega_{c}}{2\pi}\approx 6$ GHz, $\frac{g}{2\pi}\approx 450$ MHz, which gives $\frac{g}{\omega_{c}}\approx 0.075$, precisely within the coupling range of our calculations. Further we consider just two TLS’s in the cavity; the Rabi oscillation period is then $T_{Rabi}\approx 1.56\text{ ns}$. Per Fig.  1, it takes roughly 25 coherent Rabi oscillations for a fully excited initial state to be maximally enveloped by the beating frequency. Experimentally realizing this would require a polariton coherence time approximately given by

### v2

@@ I Introduction
Cavity polaritons, quasiparticles arising from the coupling of molecules and photons in an optical cavity, have become a key concept in quantum optics.[1, 2] Along with superconducting circuits, trapped ions, and cold atoms in optical lattices, these hybrid quantum platforms have lent themselves to a diverse set of applications in quantum information, materials science, and condensed matter physics. Central to their theoretical descriptions are fundamental models describing the interaction between an ensemble of two-level systems (TLS) and the cavity photons, namely, the Tavis-Cummings (TC) model, the Dicke model (DM), and the Pauli-Fierz (PF) model. Common to these models is the prediction of Rabi oscillations, which are widely recognized as the quintessential signature of quantum coherence in the light-matter interaction.[3, 4] In this letter, we predict a more subtle quantum signature that goes beyond the Rabi oscillation in revealing the fine energy structure of light-matter hybrid states (i.e., polaritons) and in differentiating between various theoretical descriptions, even in the weak coupling regime.
@@ II Model Hamiltonians
$H_{PF}=H_{DM}+\frac{g^{2}}{\omega_{c}N}\left(J_{x}\right)^{2}$ (3)
@@ III Numerical Analysis
Numerical analysis of the cavity system involves the photon counts, spectral decomposition and energy schema. Simulations were carried out for $N$ from 1 to 15, but we focus primarily on the $N=2$ resonant case for the convenience of the subsequent perturbative analysis. The parameters are $\omega_{c}=\omega_{m}=1.0$, $0.01\leq g\leq 0.12$, and an initial state consisting of a TLS ensemble in the doubly-excited symmetric state and the cavity photon in the vacuum state, $\psi(0)=\ket{s_{2},0}$. To propagate the polariton system, we combine numerical integration of the Schrödinger equation and spectral decomposition [see Eq. (4)].
@@ III.1 Average Photon Count
The average photon count, i.e., the expectation value of photon number operator $a^{{\dagger}}a\right>$, is plotted for the TC, Dicke, and PF models in Fig. 1. These plots illustrate the presence of polariton beatings in the Dicke and PF models but not in the resonant TC model. Additionally, the beating dynamics in the Dicke and PF models are similar–the only noticeable difference between the two occurs between $t=1200$ and $t=1400$, where the oscillation amplitude diminishes. The similarity can be attributed to the DSE term in the PF model, which scales with $g^{2}/N$, deeming it an order of magnitude less significant than the CRW terms. Thus, we will focus on the Dicke model in the subsequent study of the beating dynamics. As shown in SM, a similar beating pattern can be observed on the higher excitation manifold. Further, the variance of the photon count also exhibits a complex beating pattern as shown in SM, suggesting quantum features in photon statistics.[8, 9]
@@ III.3 Energy Schema & Eigenvalue Asymmetry
Evidently, the energy schema of the TC model is symmetric by construction, so $\alpha=0$ under the RWA in the TC model, but it grows with $g$ for the Dicke model, suggesting that the asymmetry $\alpha$ drives the quantum beating.
@@ V Perturbative analysis of beating
$\Delta E_{\pm}^{(0)}=\frac{\left|\braket{P_{\pm}|H_{CRW}|s_{0},0}\right|^{2}}{E_{\pm}}=\frac{g^{2}}{2(2\omega\pm\Omega)}$ (11)
@@
$\Delta E_{i}^{(4)}=\sum_{n=1}^{3}\frac{|\braket{s_{n},4-n|J_{+}a^{{\dagger}}|P_{i}}|^{2}}{E_{i}-4\omega}$ (12)
@@
We then define $\Delta E_{i}=\Delta E_{i}^{(0)}+\Delta E_{i}^{(4)}$ such that the perturbed polariton eigenenergies in the Dicke model are
@@
 $\qquad-\,8N\,\cos(\alpha t)\,\cos(\Omega\,t)\Bigg{]}$ (16)
@@ VI Experimental Relevance
Notably, modern experimental setups are capable of realizing the prediction of polariton beating. Consider a feasible physical system [17] with parameters $\frac{\omega_{c}}{2\pi}\approx 6$ GHz, $\frac{g}{2\pi}\approx 450$ MHz, which gives $\frac{g}{\omega_{c}}\approx 0.075$, precisely within the coupling range of our calculations. Further, we consider just two TLSs in the cavity; the Rabi oscillation period is then $T_{Rabi}\approx 1.56\text{ ns}$. Per Fig.  1, it takes roughly 25 coherent Rabi oscillations for a fully excited initial state to be maximally enveloped by the beating frequency. Experimentally realizing this would require a polariton coherence time approximately given by